OpenStudy (anonymous):

find dy/dx of y=e^(-5x^2)

OpenStudy (ipwnbunnies):

\[d[e^{u}] = e^{u} du\]

OpenStudy (ipwnbunnies):

I don't know if I wrote that out right.

zepdrix (zepdrix):

Ani, you remember the derivative of e^x?

OpenStudy (anonymous):

yeah its just e^x right?

zepdrix (zepdrix):

Good, same thing happening here,\[\Large\bf\sf \frac{d}{dx}e^{-5x^2}\quad=\quad e^{-5x^2}\cdot \color{royalblue}{\frac{d}{dx}(-5x^2)}\]Exponential gives us the same thing back. We just need to apply the chain rule since the exponent is more than simply `x`.

zepdrix (zepdrix):

So we need to multiply by the derivative of the exponent.

OpenStudy (anonymous):

so e^(-5x^)(-10x)

zepdrix (zepdrix):

Yayyyy good job \c:/

OpenStudy (anonymous):

lol thank you ^_^